### v1

# DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports

Abstract
As large language models (LLMs) advance, deep research systems can generate expert-level reports via multi-step reasoning and evidence-based synthesis, but evaluating such reports remains challenging. Existing benchmarks often lack systematic criteria for expert reporting, evaluations that rely heavily on LLM judges can fail to capture issues that require expert judgment, and source verification typically covers only a limited subset of explicitly cited statements rather than report-wide factual reliability. We introduce DEER, a benchmark for evaluating expert-level deep research reports. DEER comprises 50 report-writing tasks spanning 13 domains and an expert-grounded evaluation taxonomy (7 dimensions, 25 sub-dimension) operationalized into 130 fine-grained rubric items. DEER further provides task-specific expert guidance to help LLM judges assess expert-level report quality more consistently. Complementing rubric-based assessment, we propose a document-level fact-checking architecture that extracts and verifies all claims across the entire report, including both cited and uncited ones, and quantifies external-evidence quality. DEER correlates closely with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
DEER: A Comprehensive and Reliable Benchmark
for Deep-Research Expert Reports
Janghoon Han1, Heegyu Kim1, Changho Lee1, Dahm Lee1, Min Hyung Park1, Hosung Song1, Stanley Jungkyu Choi1, Moontae Lee1,2, Honglak Lee1,3 1LG AI Research, 2University of Illinois Chicago, 3University of Michigan, Ann Arbor Correspondence: hanjh9439@gmail.com
1 Introduction
Driven by the rapid advancement of large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry openai2025deepresearch; google2025gemini; anthropic2025claude; qwen3technicalreport; webthinker; manusearch; webweaver. Unlike simple information retrieval, deep research systems tackle complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results. This enables them to integrate information from diverse sources and analyze different perspectives in a balanced manner, producing reliable, evidence-based research reports deepresearch_survey; deepresearch_survey2; characterizing. These advanced capabilities enable deep research systems to achieve strong performance even on challenging benchmark tasks.
Figure 1: Deep Research System Performance Comparison. Shows the performance by type for 5 major models in the proposed benchmark.
Early evaluation of deep research systems relied on complex reasoning benchmarks gpqa; hle, which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through performance on these tasks. Subsequently, deep web search QA benchmarks were developed to directly assess systems’ web browsing and information retrieval abilities—core capabilities of deep research—by measuring multi-step search, information integration, and answer derivation browsecomp; frames; gaia; browsecomp_plus; mind2web2. More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, going beyond simple short-answer evaluation webthinker; deepbench; deepconsult; deepgym.
Despite these advancements, current methodologies for evaluating deep research systems still face key limitations in assessing expert reports. First, while prior work has proposed high-level dimensions for report evaluation, these frameworks have yet to comprehensively specify the diverse elements required for expert-level reports. In particular, many frameworks deepbench; DeepScholar-Bench; LiveResearchBench; DeepResearchArena rely on LLMs—rather than human experts—to instantiate detailed sub-dimension and concrete evaluation items, which can lead to inconsistencies and reduced reliability. Second, even with a fixed rubric, evaluations that rely on LLM judges can fail to capture issues that require expert judgment. Third, current approaches to source verification are typically limited to sections with explicit citation markers, leaving report-wide claim reliability insufficiently examined. Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
To address these limitations, we propose the Deep rEsearch Expert Report benchmark (DEER), consisting of 50 report-generation tasks spanning 13 domains, for comprehensive and reliable evaluation. DEER analyzes a wide range of expert report assessment criteria and consolidates them through expert consensus into a Deep Research Report Evaluation Taxonomy with 7 dimensions and 25 sub-dimension. Based on this taxonomy, DEER evaluates each report along two complementary components: (i) report-quality dimensions that require holistic document-level judgment and (ii) external-information dimensions that can be assessed at the claim level via fact verification. For report quality, we operationalize the taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-judge-based scoring. For external information, we introduce a fact-verification module that checks all cited and uncited claims across the report and produces quantitative measures of external-evidence quality. Finally, DEER combines these metric-based signals with rubric-based LLM evaluations to yield an integrated, multi-dimensional assessment of each expert report. Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated on DEER.
The main contributions of this study are as follows:
  • We present DEER, a systematic and interpretable benchmark for deep research, built on a hierarchical evaluation taxonomy.
  • We instantiate this taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-judge-based evaluation.
  • We propose a fact verification architecture that backtracks inter-claim dependency relations to retrieve and verify relevant citations, enabling more complete report-wide evaluation of all claims.
2 Related Works
2.1 LLM-based Evaluation Automation
With the proliferation of Large Language Models, various approaches have been proposed to automate human-centric evaluation tasks using LLMs. Most notably, the LLM-as-a-Judge method, which directly uses LLMs as evaluators, has been widely researched. JudgeLM zhu2023judgelm and G-Eval liu2023g demonstrate that models can score generated responses based on multiple criteria such as accuracy, consistency, and logic. In addition, the scope of automated LLM evaluation is expanding to pairwise comparison-based evaluations such as Chatbot Arena chiang2024chatbot, CheckList-style robustness and bias checks checkeval, and various automated scoring frameworks. However, these approaches often struggle to control bias or hallucinations in the LLM’s own judgments and do not adequately address evaluations that require complex evidence structures, such as long-form reports.
2.2 Deep Research and High-Difficulty QA Evaluation
Recently, Deep Research benchmarks have emerged that go beyond simple knowledge recall to evaluate complex reasoning, web browsing, and information integration capabilities. High-difficulty QA benchmarks such as HLE hle, GPQA gpqa, and FRAMES frames reveal the limitations of existing models by requiring long-context reasoning, expert-level thinking, and information synthesis skills. Furthermore, ManuSearch manusearch and Mind2Web-2 mind2web2 employ multi-agent web browsing and agent-as-a-judge architectures to evaluate models’ ability to replicate real-world research processes. Additionally, DeepResearchBench deepbench and WebThinker Eval webthinker present Deep Research report-centric evaluations that assess literature-based report generation and multi-document reasoning, classifying them as measuring "research capability" beyond single-answer QA. However, while these benchmarks evaluate correctness, they cannot precisely verify the validity of evidence, citations, and arguments within the report, focusing more on whether the correct answer is reached than on factuality.
2.3 Claim Verification
In automated claim verification, various approaches have been proposed to mitigate LLM hallucinations. TrustGPT huang2023trustgpt attempted to improve factuality based on RAG, and InFact rothermel-etal-2024-infact established a structural verification system by decomposing claim verification into a 6-step pipeline. Additionally, MAPLE zeng2024maple, FactDetect jafari2024robustclaimverificationfact, and ClaimCheck putta-etal-2025-claimcheck implemented sophisticated claim-level fact-checking by combining claim type classification, evidence extraction, and verification automation. Furthermore, research analyzing citation quality in detail, such as ALiiCE xu2024aliiceevaluatingpositionalfinegrained and CiteEval xu2025citeevalprincipledrivencitationevaluation, which evaluate citation accuracy and evidence alignment in academic texts, is increasing. However, most approaches focus on verifying explicit citations or single claims and fail to address issues such as implicit evidence, contextual citation inheritance, and Fair Use that appear in long-form reports.
3 Data Construction
To construct an evaluation dataset that reflects real-world usage, we analyzed 5,045 user queries collected from a deep research system and estimated the resulting topic distribution. For topic classification, we used the taxonomy of weborganizer, following prior work deepbench.
We use Humanity’s Last Exam (HLE) hle as a source of seed questions because its expert-written, high-difficulty, multi-disciplinary questions align with the expert-level topics addressed in deep research reports. To match our target topic distribution, we mapped our topic domains to the 13 HLE subject domains and sampled 50 seed questions accordingly.
Because HLE items are posed in a QA format, they require reformulation before they can serve as deep research report-generation tasks. We therefore asked domain experts (with at least a master’s degree or equivalent expertise in the relevant field) to review each item’s question, answer, and rationale to identify the underlying concepts, theories, and phenomena, and to reformulate the items as report- or paper-style topics for report generation. Each reformulated prompt was drafted by one domain expert and cross-reviewed by another expert from the same field, with iterative revisions as needed. During reformulation, we removed answer-revealing elements (e.g., factual conclusions, proofs, or specific answers) and instead provided only high-level writing directions, such as the intended scope of analysis and comparative perspectives, so that the model must develop the reasoning and narrative on its own. This reformulation shifts the task from reproducing a short answer to generating a report that requires expert-level analysis, reasoning, and interpretation. Detailed content is presented in Appendix B.
4 Approach
4.1 Overview
Figure 2: Overview of the DEER evaluation framework. (a) Research question and expert guidance generation from real-world deep research queries. (b) Construction of the Deep Research Evaluation Taxonomy consisting of 7 dimensions, 26 criteria, and 130 granular rubrics. (c) The DEER evaluation pipeline, integrating expert-guided LLM-as-a-judge scoring with claim extraction and information verification to assess deep research reports.
To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis characterizing; deepresearch_survey; deepresearch_survey2. Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy with 7 major dimensions (§ 4.2), which we group into 5 report-quality dimensions and 2 external-information dimensions. Report-quality dimensions primarily assess synthesis and presentation via holistic, document-level judgment, whereas external-information dimensions assess external evidence acquisition and use via claim-level verification.
Based on these differences in evaluation characteristics, we propose a hybrid architecture that integrates methodologies tailored to each component, as illustrated in Figure 2. Expert-Guided Report Quality Assessment (§ 4.3) uses an LLM-as-a-judge approach, combining a fixed set of granular evaluation items with report-specific Expert Evaluation Guidance authored by domain experts, to assess report quality against expert-designed criteria. Information Verification (§ 4.4) performs metric-based evaluation by automatically extracting claims and citations and verifying claims against external sources, producing quantitative metrics of external evidence sufficiency and integrity. Scores from the rubric items and verification metrics are aggregated into 7 dimension scores, which are then combined into an overall report score.
Major Dimensions Sub-criteria
Request Completeness Request Completeness, Scope Boundary, Value Validity
Evidence Validity Numeric Accuracy, Logical Support
Structure Consistency Structure Intro, Structure Body, Structure Conclusion, Structure Section
Narration Style Report Form, Writing Quality, Paragraph, Reader Friendliness
Ethics Compliance Sensitive Issues, Safety Impact, Perspective Balance
Information Sufficiency Source Support, Information, Citations, References
Information Integrity Claim Factuality, Citation Validity, Reference Accuracy, Reference Quality, Reference Diversity
Table 1: Deep Research Report Evaluation Taxonomy: 7 Major Dimensions and 25 Sub-criteria
4.2 Deep Research Report Evaluation Taxonomy
In the absence of systematic evaluation criteria for deep research reports, we construct an evaluation taxonomy by synthesizing expert report assessment standards from multiple fields. We draw on internationally established standards spanning 20 domains of expertise, including systematic research reporting guidelines, technical and professional report writing and evaluation guidelines, and academic publishing norms, for a total of 80 standards. A panel of experts with experience in deep research system development and academic reviewing iteratively analyzes and consolidates these standards to derive report-quality dimensions. We then introduce complementary external-information dimensions that reflect deep research–specific characteristics of external evidence use and verification, resulting in 7 major dimensions and 25 detailed criteria in total (Table 1). We validate the taxonomy through independent review by 10 experts from diverse domains. This hierarchical taxonomy organizes diverse quality elements into structured evaluation dimensions and sub-dimension, enabling systematic, multi-dimensional evaluation. It facilitates diagnosis of report strengths and weaknesses and guides systematic improvement. Full descriptions, validation procedures, and the mapping from source standards to our criteria are provided in Appendix C.
4.3 Expert-Guided Report Quality Assessment
Existing evaluation methods have two limitations. First, existing methods apply broad, high-level criteria to long-form reports, giving LLM judges substantial discretion over what to focus on. As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) even focus on non-critical or superficial sub-aspects, increasing variance across judges and reducing evaluation consistency webthinker; deepconsult; deepgym. Second, even with a clear evaluation focus, assessing correctness and completeness often requires domain knowledge, which LLM judges may lack. As a result, they may miss subtle but important issues such as non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies deepbench. This study addresses these issues by (1) making the evaluation well-specified through a fixed set of fine-grained rubric items and (2) providing task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, helping LLM judges identify hard-to-detect errors and omissions.
Granular Rubric Design. Broad evaluation criteria can lead to inconsistent LLM-judge-based scoring. Recent work addresses this by decomposing broad criteria into granular rubrics checkeval; rocketeval; expertlongbench. For deep research report evaluation, deepbench follows a similar direction by using an LLM to generate task-specific evaluation criteria. While this narrows the evaluation focus, it raises two concerns: whether the generated criteria are sufficiently complete and capture what matters most in expert reports, and whether dynamically changing criteria enable interpretable, comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation. Building on the taxonomy in Table 1, a panel of experts operationalizes the 25 sub-dimensions into 40 evaluation criteria and translates each criterion into concrete, checkable rubric items. Rubric items are organized into two aspects: coverage (whether the required components are present at all relevant points in the report and addressed in full whenever they occur) and quality (the degree to which the targeted components are executed to a high standard in the report). This process yields 130 scorable rubric items in total. [^1] The 40 criteria were validated through independent review by 10 experts following the procedure in § 4.2. The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix E for the rubric structure and examples).
Expert Evaluation Guidance. Even with a well-specified, granular rubric, scoring individual items can still require judgment informed by deep domain knowledge, especially for specialized and complex tasks—for example, to assess whether the report’s reasoning is sound, its calculations are correct, and its coverage and depth are adequate for the topic. In such cases, evaluators may miss subtle domain-specific errors or omissions that human experts would catch. To mitigate this, we provide task-specific expert evaluation guidance that enumerates the required content elements and expert expectations for each task, presented as concrete, checkable statements that evaluators can verify.
For each task, an expert with at least a master’s degree in the relevant field was assigned to detail the key concepts, theories, data, and analytical perspectives that the report must cover from an expert perspective. The expert first modified the original query to naturally include the essential content of the topic (Section 3), and then wrote the Expert Evaluation Guidance based on the modified query, specifying the content elements that must be included. This approach is based on the logic that if a report responds appropriately to the modified query, it should naturally contain the key content of the topic, so the omission of key content can be clearly judged as a quality issue. Each written Guidance was verified through cross-review by another expert in the same field. See Appendix B for detailed methodology.
4.4 Information Verification
To automate fact verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations. In reality, many sentences implicitly inherit preceding evidence or construct new claims through contextual connections. Excluding these implicit claims from verification makes the reliability assessment of the entire document incomplete. Our Information Verification Module proposes an architecture that extends the verification scope to the entire document by tracking sentence-level logical flows.
Implicit Claim & Back-Tracking
This module analyzes contextual correlations to convert even uncited sentences into verifiable units. Specifically, when sentence $s_{i}$ implicitly cites the content of a previous sentence or paragraph, we use a Back-Tracking mechanism to trace the reference source. To this end, the LLM identifies the set of preceding sentences $R(s_{i})$ that $s_{i}$ depends on, and the valid citation set $\mathcal{V}(s_{i})$ used for actual verification is defined as follows:
$\mathcal{V}(s_{i})=\mathcal{C}(s_{i})\cup\bigcup_{k\in R(s_{i})}\mathcal{C}(s_{k})$ (1)
where $\mathcal{C}(s_{i})$ is the explicit citation of $s_{i}$. This structure allows verification by inheriting citations from referenced preceding sentences even if the sentence itself has no citation ($\mathcal{C}(s_{i})=\emptyset$). This is a key element that substantially expands the verification scope in policy and research reports containing many implicit claims.
Batch Processing for Claim Extraction
When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle problem, which leads to the omission of important claims liu2023lostmiddlelanguagemodels. To address this, we apply a Batch Extraction strategy that maintains the context of the entire document $D$ while parallel-processing only the target sentences divided into small batches $B_{j}$:
$\mathcal{A}=\bigcup_{j=1}^{M}E(D,B_{j}),\quad D=\bigcup_{j=1}^{M}B_{j}$ (2)
This increases the model’s focus to prevent missing important claims, and experimental results showed a significant improvement in extraction recall compared to the single processing ($E(D,D)$) method. Although the number of API calls increases, we confirmed that this is an essential strategy for securing high coverage in reports dealing with hundreds of claims.
Scalable Verification Architecture
The extracted claims are passed to the fact verification stage, where the biggest technical challenge is the trade-off between accuracy and cost. Verifying every claim individually yields high accuracy, but cost increases linearly with long documents. To address this, we introduce two scaling strategies.
First, we apply Grouped Claim Verification, which bundles multiple claims citing the same source into a single input for verification. Experimental results showed that increasing Group Size from 1 to 10 raised accuracy from 0.78 to 0.89 and reduced cost per claim by 85% (Section 6.5). This indicates that grouped input helps in understanding the overall structure of the document and verifying claims, while reducing verification overhead. Second, greater API cost reduction is possible through the Context Retrieval strategy, which selectively provides only the context necessary for fact verification. This method achieved an API cost reduction of approximately 24 times, but a 6% drop in accuracy was observed, indicating a clear trade-off.
This hierarchical pipeline (Batch Extraction $\rightarrow$ Grouped Verification $\rightarrow$ Context Retrieval) is designed to achieve both high-precision verification and cost efficiency even in long reports containing hundreds of claims.
5 Experiment Setup
5.1 Implementation Details.
We use the LLM-as-a-judge method to evaluate report quality. The evaluation model takes as input the task query, the report to be evaluated, expert guidance, and detailed rubrics and evaluation criteria. The evaluation is conducted across 5 dimensions: Request Completeness, Evidence Validity, Structure Consistency, Narration Style, and Ethics Compliance, with each rubric item scored on a 1–10 scale. The model generates a score and a short rationale justifying that score for each item, and the final output is in JSON format consisting of (rationale, score) pairs for each item.
Information Integrity and Information Sufficiency are evaluated through a separate verification pipeline, the Information Verification Module. This module extracts claims from the report and assesses their factuality by comparing them with external evidence obtained via web searches. The verification results are calculated as detailed metrics such as Claim Factuality and Citation Validity.
The evaluation is hierarchical, computing the final quality score by sequentially aggregating the averages of lower-level rubric items to higher levels, using the scores collected above.
Model Req. Comp. Evid. Valid. Struct. Cons. Narr. Style Info. Int. Info. Suff. Ethics Mean
General LLM
qwen3-235b-fast 6.10 6.59 9.04 8.66 3.30 4.40 8.58 6.67
gemini-2.5-pro-fast 6.08 7.02 9.29 8.87 2.92 4.43 8.89 6.79
claude_opus4.1_fast 6.04 6.47 8.96 8.95 4.76 4.81 8.78 6.97
gpt5_fast 5.61 6.63 8.80 8.46 2.18 3.45 8.82 6.28
LLM+Thinking
qwen3-235b-think 6.77 6.53 9.32 8.92 3.28 4.39 8.75 6.85
gemini-2.5-pro-think 6.41 7.66 9.55 9.20 3.86 4.97 9.37 7.20
claude_opus4.1_think 5.93 6.41 9.05 9.04 4.72 5.12 8.99 7.04
gpt5_think 7.69 8.37 9.57 9.32 4.19 5.42 9.37 7.70
LLM+Thinking+Search
qwen3-235b-think_search 6.43 6.31 9.07 8.82 3.60 4.49 8.44 6.74
gemini-2.5-pro-think_search 6.61 7.58 9.50 9.11 4.62 5.02 8.97 7.34
claude_opus4.1_think_search 6.03 6.62 9.04 9.03 5.17 4.80 8.85 7.08
gpt5_think_search 7.88 8.26 9.39 9.04 6.42 6.44 9.38 8.12
Deep Research
webthinker 5.27 5.76 8.48 8.62 7.57 6.92 8.68 7.33
qwen3-235b-deep 4.95 6.00 7.28 8.31 7.13 8.29 8.66 7.23
gemini-2.5-pro_deep 6.08 6.81 8.81 8.78 7.13 7.87 8.62 7.73
claude_opus4.1_deep 6.60 6.91 8.90 8.85 6.79 7.66 8.75 7.78
gpt5_deep 6.26 7.11 9.20 8.93 7.64 7.13 8.99 7.89
Table 2: Baseline models’ expert report evaluation results. In each column, the highest score is marked in bold and the second highest score is underlined.
5.2 Baseline Models
In this study, to compare the expert report generation capability of LLM-based systems, we established four categories of baseline models. Each category is designed to increase in complexity incrementally, ranging from simple language model response generation to specialized research systems that combine search and deep reasoning.
(1) General LLMs This is the basic LLM setting that generates responses in a single pass (inference), serving as a baseline for measuring response quality in language-generation-based models. This category includes qwen3-235b-fast qwen3technicalreport, gemini-2.5-pro-fast google2025gemini, claude_opus4.1_fast anthropic2025claude, and gpt5_fast openai2025gpt5system models.
(2) LLMs with Reasoning This setting uses the reasoning function on the same LLMs as in (1), extending the internal reasoning step before answer generation. This category includes qwen3-235b-think, gemini-2.5-pro-think, claude_opus4.1_think, and gpt5_think models.
(3) LLMs with Reasoning + Web Search This configuration combines the LLMs with the Reasoning setting and web search, evaluating scenarios in which external information is used during reasoning. This category includes qwen3-235b-think_search, gemini-2.5-pro-think_search, claude_opus4.1_think_search, and gpt5_think_search models.
(4) Deep Research Frameworks (Open-Source & Commercial) This is the upper baseline, comprising specialized deep research frameworks and commercial research support systems. We used webthinker webthinker as the open-source framework, and commercial systems include qwen3-235b-deep, gemini-2.5-pro_deep, claude_opus4.1_deep, and gpt5_deep. Detailed configurations and hyperparameters for all baseline models are summarized in Appendix A.
6 Experiments
6.1 Main Results
Summarizing Table 2, all models show generally high performance in metrics related to form, style, and ethics, such as Structure Consistency (Struct. Cons.), Narration Style (Narr. Style), and Ethics Compliance (Ethics). However, they consistently show relatively lower scores on metrics that assess how faithfully they address the requirements in the user’s query and provide persuasive evidence in terms of calculation, methodology, and logical development, such as Request Completeness (Req. Comp.) and Evidence Validity (Evid. Valid.). This indicates that while the latest LLMs are already quite proficient in structuring reports, writing naturally, and adhering to basic ethical standards, their ability to accurately meet user intent at an expert level and support it with sufficiently robust argumentation and evidence structures remains limited. Reasoning models showed consistent improvements across almost all evaluation metrics relative to General LLMs, suggesting that simply expanding the reasoning budget without additional tools improves coverage of requirements, logical development, evidence selection, and narrative consistency. Models combining this with web search show additional gains in information-related metrics. The Deep Research category consistently achieves higher scores in Information Integrity (Info. Int.) and Information Sufficiency (Info. Suff.) than other categories, providing the strongest baseline for the broad collection and organization of external documents.
However, it is characteristic that the Deep Research category consistently records lower scores than reasoning-based single LLMs in the remaining key evaluation axes, especially Req. Comp. and Evid. Valid. These results suggest that the large amount of external material injected through search in the Deep Research configuration may cause the reference point for report development to gradually shift from the query itself to external knowledge, inducing a kind of retrieval-induced drift in which information rather than blurs the problem definition and argumentation structure. Therefore, this demonstrates that simply increasing the volume of search and citations does not guarantee the quality of expert reports.
6.2 Fine-grained Analysis of Evaluation Results
Figure 3: Heatmap visualizations of expert report evaluation results. (a) Criteria-wise scores across detailed evaluation categories. (b) Domain-wise scores averaged over two representative tasks randomly sampled from each domain.
Figure 3(a) decomposes the upper evaluation dimensions into detailed items to summarize the performance patterns of current Deep Research and LLM-based systems in expert reports. In Request Completeness, scores across sub-items are generally low, with scope the lowest, suggesting a lack of clarity about what the report covers and excludes, and under what assumptions and limitations the discussion is developed. In Evidence Validity, the score for the numeric item is slightly lower than that for logical, indicating that current systems make mistakes relatively more frequently in terms of quantitative accuracy, such as calculation processes, used formulas/statistical models, and selection of indicators/units, rather than logical development itself. In Information-related detailed items, ref_diversity and refs_amount appear commonly low, meaning that even if the overall structure and narrative are compliant, the number and diversity of sources actually referenced are limited.
Figure 3(b) shows the domain-specific performance of Deep Research agents as the average of two randomly selected tasks in each domain (domain-specific average scores for all tasks are presented in Appendix LABEL:app:domain-results). While most domains are distributed in a relatively similar range around 7 points, in the case of Physics, most Deep Research agents except gpt5_deep consistently record the lowest scores compared to other domains, and relatively large deviations between Deep Research agents also appear in CS, History, and Linguistics.
6.3 Correlation with Human Evaluation
To verify the validity of the proposed evaluation method and analyze the contribution of each component, we conducted experiments to assess its correlation with human expert evaluation. For 15 tasks across 5 domains, 3 reports randomly selected from those generated by the Deep Research systems used in 6.1 were used as evaluation targets. Each report was evaluated by two independent experts, collecting a total of 90 human evaluation samples. To assess the alignment between the evaluation method and human experts, we used the Pearson correlation coefficient, the Spearman rank correlation coefficient, and pairwise agreement deepbench. Pairwise agreement is a metric that measures the proportion of cases in which our evaluator chose the same priority as the human expert when comparing two reports, indicating how consistently it reproduces human relative judgments across tasks. Evaluations were performed on the same reports using 5 LLM-based evaluation models (GPT-5, GPT-5-mini, Claude Opus 4.1, Claude Sonnet 4.5, Gemini 2.0 Pro), and the correlation between the average of the 5 model evaluations and the human evaluation was analyzed.
Evaluation Method Pearson Spearman Pairwise
$r$ $\rho$ Agr.
Vanilla 0.64 0.61 0.66
+ Dimensions 0.67 0.65 0.80
+ Granular Checklist 0.53 0.46 0.72
+ Expert Guidance 0.75 0.71 0.84
Inter-Human 0.81 0.74 0.79
Table 3: Average human correlation across five AI models with incremental addition of evaluation components. Inter-Human shows inter-annotator agreement.
Table 3 shows how the correlation with human evaluation changes as evaluation components are added step by step. Starting from the Vanilla deepbench setting, which judges overall quality without explicit evaluation criteria, we sequentially expanded by introducing evaluation dimensions (+Dimensions), decomposing them into granular rubrics (+Granular Rubric), and finally adding task-specific expert guidance (+Expert Guidance) to measure the impact of each component. The Vanilla setting showed significant correlation with human experts even without explicit criteria, but this correlation decreased as evaluation dimensions (+Dimensions) were introduced and decomposed into detailed items (+Granular Rubric). Since report evaluation is a high-dimensional task that requires simultaneous consideration of various elements, such as logic, concepts, analysis, and domain knowledge, across long and complex texts, each rubric item comprises broad, interpretive judgment units rather than narrow rules. It is analyzed that the increase in the number of items led to an increase in interpretive variability rather than refinement, resulting in a drop in human correlation. Adding task-specific expert guidance (+Expert Guidance) clarifies the interpretation method for each criterion, restoring consistency and significantly improving correlation across all metrics. Writing longer rationales (+Longer Rationale) further improves correlation by enriching the basis for each item. In conclusion, DEER tends to achieve relatively stable criterion application and high alignment with human experts, even in report evaluation settings that must consider multiple evaluation dimensions simultaneously.
6.4 Inter-evaluator Agreement (IEA)
To verify that the proposed evaluation method yields consistent results across different LLMs, we measured inter-evaluator reliability among the 5 evaluation models used previously IAA; checkeval. We measured the agreement between models using Krippendorff’s $\alpha$ and ICC (Intraclass Correlation Coefficient). High agreement suggests that the proposed method possesses versatility and stability that are not dependent on a specific LLM.
Method Krip. $\alpha$ ICC(2,1) ICC(2,k)
Vanilla 0.46 0.48 0.82
+ Dimensions 0.32 0.37 0.75
+ Granular Checklist 0.17 0.24 0.62
+ Expert Guidance 0.55 0.56 0.87
Table 4: Inter-evaluator reliability across five LLM-based evaluation models. Krip. $\alpha$: Krippendorff’s alpha. Higher values indicate more consistent evaluations.
Table 4 shows how the agreement between evaluation models changes as evaluation components are added step by step. The Vanilla setting showed an appropriate level of inter-model agreement, but agreement decreased significantly when introducing evaluation dimensions (+Dimensions) and decomposing them into detailed items (+Granular Rubric). As explained earlier, this is because rubric items consist of interpretive judgment units, so the increase in the number of items increased the variability in interpretation across models. Adding expert guidance (+Expert Guidance) significantly improved agreement across all metrics, restoring evaluation consistency. In particular, the variation focusing only on the top rubric items achieved the highest agreement. This shows that expert guidance clarifies the interpretation of evaluation criteria, enabling consistent evaluation even between different LLM evaluation models.
6.5 Information Verification Module Evaluation
In this section, we evaluate the proposed Information Verification Module by dividing it into two stages: claim extraction and fact verification.  [^2] Detailed content is covered in Appendix F.
Claim Extraction Analysis
We compared and analyzed 728 claims identified by human experts (Ground Truth) and LLM extraction results using a Paragraph-Level LLM Judge (Table 5). As a result of measuring semantic Recall beyond simple count comparison, GPT-5-mini showed significantly higher performance compared to GPT-4.1. In particular, the GPT-5-mini (Batch 5) setting recorded a Recall of 0.9333 for Type A, B, and C claims requiring external evidence, capturing human expert claims almost perfectly. Also, the number of claims extracted per paragraph (Claims/Para) was the highest at 6.95, confirming that detailed information in the document is extracted densely without omission. On the other hand, GPT-4.1 showed relatively lower Recall (about 0.83) and density (about 4.8), suggesting a tendency for the model to excessively summarize information or omit some details. In conclusion, for precise fact-checking, a strategy of maximizing recall by operating high-performance models like GPT-5-mini with small batch sizes (Batch Size 5, 10) is essential.
Table 5: Comparison of Claim Extraction Performance across Models and Batch Sizes. Recall is measured at the Paragraph-Level using LLM Judge.
Model Batch Size Recall Claims/Para
Ground Truth - - 7.22
GPT-4.1 5 0.9057 5.00
10 0.8897 4.84
20 0.8897 4.65
GPT-5-mini 5 0.9333 6.95
10 0.9264 6.22
20 0.9218 6.05
Claim Verification Performance
In the claim verification stage, we conducted an ablation study on grouping and retrieval settings using GPT-4.1 (Table 6). When applying Grouped Claim Verification without retrieval, the F1 score remained comparable (91 → 90), while the cost per claim decreased by approximately 82%, demonstrating that batching claims provides substantial efficiency gains with minimal performance loss.
Additionally, enabling retrieval further reduced the cost to \$0.004 per claim but lowered the F1 score to 85. Therefore, for scenarios requiring high reliability, grouped verification without a retrieval configuration offers the best balance, while with a retrieval setting, it is preferable when maximizing cost efficiency is the primary objective.
Grouped Retrieval F1 Cost (\$/Claim)
0.91 0.083
20 0.90 0.015
20 0.85 0.004
Table 6: Ablation study on GPT-4.1. Grouping significantly improves cost efficiency with minimal F1 degradation; retrieval further reduces cost at the expense of accuracy.
7 Conclusion
We introduce DEER, a comprehensive benchmark for evaluating Deep Research agents, featuring a multi-dimensional taxonomy and a hybrid evaluation framework that combines LLM-based qualitative assessment with a retrieval-augmented Information Verification Module. Our experiments reveal that while state-of-the-art models excel in structure and style, they still lag behind human experts in critical areas like evidence validity and information sufficiency, although "Deep Research" systems show promising improvements. By demonstrating a high correlation with human judgments, DEER offers a scalable standard for advancing autonomous research agents toward trustworthy, expert-level performance.
Limitations
While our evaluation framework relies on LLM-based judges, which may inherently exhibit biases relative to human experts, our extensive validation shows a high correlation with human judgment, suggesting that these biases are systematic and manageable. Furthermore, although our current benchmark focuses on text-based reports, this specialization enables a deeper, more rigorous analysis of information integrity and logical coherence, laying a solid foundation for future extensions to multimodal research tasks.
Acknowledgments
Benchmark Human rubrics Expert curated Open- ended Non-tech domains LLM judge Claim fact-check Inter- pretability Systematic criteria Avg. rubrics
AcademicBrowse AcademicBrowse
BrowseComp browsecomp
Mind2Web2 mind2web2 $\triangle$ 50
ResearchBench ResearchBench
ResearcherBench ResearcherBench $\triangle$ $\triangle$ 14
LiveDRBench characterizing $\triangle$
SPOT SPOT $\triangle$
ExpertLongBench expertlongbench 16
ReportBench ReportBench $\triangle$
DeepScholar-Bench DeepScholar-Bench $\triangle$ $\triangle$
DeepResearchBench deepbench $\triangle$ $\triangle$ 25
LiveResearchBench LiveResearchBench $\triangle$ $\triangle$
DeepResearchArena DeepResearchArena $\triangle$ $\triangle$
DeepResearchGym deepgym $\triangle$
ResearchRubrics ResearchRubrics $\triangle$ 26
DREX (Ours) 130+
Table 7: Comparison of DREX with representative deep research benchmarks. Here, indicates full support, no support, and $\triangle$ partial support. For Claim fact-check, $\triangle$ means that only a subset of claims (e.g., explicitly cited or gold-labeled ones) are checked rather than all explicit and implicit claims. For Interpretability, $\triangle$ means that the benchmark provides only coarse, dimension-level insight (e.g., a few high-level scores) without shared checklist-level diagnostics across tasks.
Appendix A Comparison with Deep Research Benchmarks
In this section, we compare existing deep research benchmarks with our methodology. Table 7 extends and modifies the comparison table proposed in ResearchRubrics ResearchRubrics, and adds new evaluation axes to summarize how DREX differs from prior work. The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used). On top of this, we add three axes that we consider essential from the perspective of deep research report evaluation: (1) Claim fact-check, (2) Interpretability, and (3) Systematic criteria. Systematic criteria indicate whether the evaluation criteria form a shared hierarchical ontology across tasks, grounded in external expert norms and guidelines.
The first block of the table consists of benchmarks for evaluating search and browsing agents. AcademicBrowse AcademicBrowse, BrowseComp browsecomp, and Mind2Web2 mind2web2 evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources. These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research. ResearchBench ResearchBench and ResearcherBench ResearcherBench evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics. Characterizing Deep Research characterizing and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning. SPOT SPOT measures how well a system can detect critical errors in published papers, and ExpertLongBench expertlongbench evaluates the ability to generate long-form expert text without external search. These benchmarks finely probe partial abilities such as hypothesis generation, claim discovery, error detection, and expert writing, but do not directly evaluate the quality of full reports produced by web- or literature-based deep research agents.
The third block targets benchmarks that evaluate the quality of deep research reports themselves. ReportBench ReportBench and DeepScholar-Bench DeepScholar-Bench measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations. DeepResearchBench deepbench and LiveResearchBench LiveResearchBench evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy. DeepResearchArena DeepResearchArena automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE). DeepResearchGym deepgym provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality. ResearchRubrics ResearchRubrics provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DREX in terms of evaluation setting. DeepResearchBench and LiveResearchBench are most similar to our web-based deep research report evaluation setting, but both rely heavily on LLMs to define evaluation criteria and assign scores, and their fact-checking is applied only to a limited subset of claims with explicit citations. ResearchRubrics has the advantage that domain experts write 20–40 rubric items for each deep research prompt, enabling fine-grained capture of what constitutes a “good” answer, but it does not perform source-level claim fact-checking, and the rubric structure differs across prompts, making it challenging to interpret system behavior consistently along shared dimensions and rubrics. In contrast, DREX extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 130 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance. This allows DREX to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
Table 8: Comparison of Information Verification Pipelines.
Benchmark Explicit Verif. Implicit Verif. Global Context Key Limitation vs. DREX
DeepResearch Bench
deepbench
Ignores uncited sentences; fails to detect hallucinations in transitional logic.
DeepResearchGym
deepgym
Limited to explicitly cited claims; only calculates citation precision/recall metrics.
ResearcherBench
ResearcherBench
Treats uncited claims simply as “ungrounded” (empty URL) without attempting context-based verification.
ReportBench
ReportBench
$\triangle$ Verifies non-cited claims via external web search voting, ignoring internal document grounding.
DeepScholar-Bench
DeepScholar-Bench
$\triangle$ Relies on physical distance (sliding window $w$); misses long-range semantic dependencies.
DeepResearch Arena
DeepResearchArena
$\triangle$ Evaluates Source $\to$ Report coverage (summarization) rather than Report $\to$ Source verification; misses uncited hallucinations.
LiveDRBench
characterizing
Relies on matching against static Ground Truth claims; lacks dynamic verification of citations against web sources.
SPOT
SPOT
Evaluates internal error detection against static ground truth; penalizes valid but unannotated error predictions (false positives).
[…]
Appendix B Data Construction Details
B.1 Topic Domain Analysis
To configure Deep Research tasks (queries), we first analyzed 5045 in-house user queries collected from Deep Research services to ensure a balanced distribution of tasks across domains. The domain distribution of these queries is presented in Figure 4, from which we finalized 8 topic domains, including the top 7 areas and an "Others" category that groups the rest.
B.2 HLE Subject Mapping
Although actual Deep Research service logs are collected in a research-oriented usage context unlike general QA systems, since the majority of users are not limited to experts in specific academic fields, individual queries are not designed with the premise of generating expert-level academic reports or papers immediately. In addition, they include partial queries that depend on the ongoing session context or practical/pragmatic information needs, so using them as-is for evaluation research tasks may lead to a mismatch with the high-difficulty research tasks we aim for in scope and format.
Accordingly, in this study, while maintaining the 8 topic domain composition derived from actual Deep Research logs, we defined subjects from Humanity’s Last Exam (HLE) hle that semantically correspond to each domain and used HLE items belonging to those subjects as seed tasks for each domain. At this time, only items within the difficulty range currently processable by LLMs among all HLE items were selected through a preliminary performance evaluation, and the selected items were remapped to the previously defined Deep Research topic domains to obtain a higher-difficulty academic task composition while preserving the actual usage distribution. The correspondence between each topic domain of the actual Deep Research service and HLE subjects is summarized in Table 9.
B.3 Conversion from HLE QA to DeepResearch Reports
Since the selected HLE items are initially configured in QA format, they are not suitable for deep report generation tasks as they stand. In this study, experts reviewed the question-answer-rationale of each item to identify the core concepts, theories, and phenomena covered, and then reconstructed them into research-oriented task queries suitable for the Deep Research usage context. At this time, deterministic elements such as specific answers, factual conclusions, or proofs presented in HLE were removed from the prompt, and instead, only high-level writing directions such as analysis scope, comparative perspectives, and issues to be discussed were presented to induce the model to develop the logical structure and conclusion on its own.
In addition, if necessary, the basic structure of the output, such as "Introduction–Related Work–Method–Analysis–Discussion," was briefly specified to make the task more similar to the long-form research report generation expected in actual Deep Research use cases. Through this conversion process, simple answer-type QA items were transformed into a report-generation task set with HLE-level high-quality and high-difficulty problem structures, while following the actual Deep Research usage domain distribution.
Figure 4: Topic domains extracted from real-world Deep Research service logs
In-house Domain Humanity’s Last Exam Subject
Finance & Business Economics
Software Development Computer Science, AI
Science & Technology Mathematics, Physics, Chemistry
Industrial / Hardware Engineering
Education & Jobs Education
Health Biology, Psychology
Others History, Linguistics, Philosophy
Table 9: Mapping between in-house topic domains and Humanity’s Last Exam subjects. Combines "Software Development" and "Software" categories.
Example Research Task (MDP / Value Iteration) Write a research report that explores the conditions under which value iteration in Markov Decision Processes converges geometrically to the optimal value function. Your report should provide precise definitions of the key concepts involved, examine the roles of rewards and discounting in ensuring convergence, and analyze both theoretical and practical factors that influence the algorithm’s behavior. Discuss circumstances where convergence occurs, where it may fail, and highlight open questions or limitations. The report should be structured as a rigorous technical investigation, integrating mathematical reasoning, conceptual explanations, and illustrative examples to support the analysis.
Appendix C Deep Research Report Evaluation Taxonomy Details
C.1 Dimension and Criteria Specification
In this section, we detail the 7 dimensions and 26 detailed criteria of the Deep Research Report Evaluation Taxonomy presented in the main text.
Request Completeness
This dimension assesses how systematically and comprehensively the report addresses the user query with respect to the required elements, analysis scope, and final key message. Request Completeness assesses whether all requirements presented in the query are included without omission and developed with sufficient depth. Scope Boundary assesses whether the covered content, excluded items, assumptions, and limitations are clearly defined and consistently maintained. Value Validity assesses whether the final key message encompasses the entire query, is sufficiently specific and practical, and is logically consistent with the evidence and analysis results presented in the report body.
Evidence Validity
This dimension evaluates how accurate and valid the report’s figures and arguments are in terms of calculation, methodology, and logical development. Numeric Accuracy examines whether calculation processes, used formulas/statistical models, and indicators/units are presented without error, are appropriate for the problem context, and are expressed transparently enough for a third party to reproduce and verify. Logical Support evaluates whether the argument is developed consistently with the topic, necessary background/assumptions/inference steps are specified, and major claims and counter-argument responses are persuasively supported without leaps based on facts, data, and interpretations.
Structure Consistency
This dimension evaluates whether the introduction, body, and conclusion, and the structure of each section of the report, are organized consistently with the topic and scope. Introduction checks whether it concisely and clearly presents the topic, problem, significance, and basic scope. The body checks whether the step-by-step argument is developed without omission or deviation, in accordance with the structure and scope presented in the introduction. Conclusion checks whether it synthesizes the body content to complete the purpose of the introduction without introducing new claims or evidence. Section-Level checks whether each section supports the overall structure through clear organizational principles and appropriate connections.
Narration Style
This dimension assesses whether the report faithfully follows the required format and style and is expressed in a way that the reader can read and understand the content without difficulty. The report form checks whether external requirements, such as document length and section system, conform to professional report practices. Writing Quality checks whether sentences are concise and accurate, and whether terminology and tone are consistent. Paragraph checks whether paragraphs are sufficiently developed around a single point and naturally integrated with structural auxiliary elements. Reader Friendliness assesses whether subheadings and simple explanations/examples are used to guide the reader in following complex content.
Ethics Compliance
This dimension assesses whether the report is written ethically and responsibly regarding sensitive issues, potential harm, and perspective balance. Sensitive Issues checks whether sensitive topics such as politics, race, and gender are addressed with neutral, fair language and a balanced perspective. Safety & Impact checks whether negative impacts, side effects, and potential misuse of proposals/technologies/research results are adequately reviewed, and whether specific method presentations are dangerous. Perspective Balance assesses whether the discussion is balanced by appropriately including related perspectives and opposing views without bias toward a specific position.
Information Sufficiency
This dimension assesses whether the information required to answer the request has been adequately secured and presented in terms of quantity and scope. Source Support evaluates whether verifiable evidence or reliable sources are provided to support claims that require external evidence. Information evaluates whether reliable facts and claims are sufficiently presented. Citations evaluate whether sufficient citations are made at necessary points. References are evaluated to determine whether the number of actually used references reaches an appropriate level.
Information Integrity
This dimension assesses the factuality of the external information used in the report and the reliability and diversity of the citations/sources supporting it. Claim Factuality measures the proportion of verifiable claims that are determined to be factual. Citation Validity measures the proportion of citations that actually support the claim among citations attached to each claim. Reference Accuracy measures the proportion of sources that correctly support the argument among the presented/used references. Reference Quality assesses whether the sources used are reproducible and reliable. Reference Diversity assesses whether the evidence maintains a sufficient level of source diversity without becoming overly concentrated on a few documents.
C.2 Construction Procedure and Evidence Mapping
To systematize and standardize universal and essential elements of expert report evaluation, this study synthesized and normalized evaluation standards across the natural sciences, engineering, and social sciences. To this end, we analyzed authoritative standards and guidelines widely used for writing and evaluating expert reports, such as guidelines for writing/reviewing academic papers and research reports, guidelines for systematic reviews/literature reviews, academic publication norms, and guidelines for writing/evaluating policy/regulatory/market analysis reports and consulting/advisory reports. We extracted the common required elements from these materials and integrated/generalized overlapping or domain-specific items.
To verify the validity of the drafted dimensions and detailed criteria, this study conducted cross-validation with 10 independent experts representing 6 domains (Computer Science, Artificial Intelligence, Biology, Chemistry, Mathematics, Psychology). Two experts per domain reviewed the suitability of each item, focusing on factors considered necessary in light of domain practices, such as whether each dimension and criterion is actually a meaningful and necessary evaluation standard in that domain, whether the definition and scope of application are excessively ambiguous, and whether it unnecessarily overlaps with other criteria. Based on this, they made binary (pass/fail) evaluations for each criterion. The final evaluation framework included only the criteria that passed this cross-validation process.
The following presents the mapping of which external standards were referenced for each evaluation dimension.
Appendix D Foundations of Deep Research Report Evaluation Taxonomy
Request Completeness
Request Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian2021PRISMA; Boutron2010CONSORT; ISO25010; WWC_Handbook; Percie2020ARRIVE; GRI_Standards; STROBE_2007. Scope Boundary reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian2021PRISMA; NIST_AI_RMF; AHA2024Standards; AEA2024Data; OSF_Prereg; NASA_SE_Handbook; SEC_Reg_SK; SPIRIT_2013. Value Validity corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt2013GRADE; APA_JARS2025; NBER2024; IES_SEER; OECD_Policy_Eval.
Evidence Validity
Numeric Accuracy reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt2013GRADE; JCGM_GUM; ACM_Badging; IUPAC_Green; IUPAP_Red; Econometric_Soc; AERA_APA_NCME_Standards; CODATA_2022; OECD_Quality; IFRS_Standards; ISO30414; IPCC_2019. Logical Support aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor EMS_Code2025; APA_GoodPractices; AERA2006Standards; AMS_Handbook; Fed_Rules_Evidence; COREQ_2007.
Structure Consistency
Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; IFRS_Integrated; CARE_2013. Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; CF_Conventions; CARE_2013. Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian2021PRISMA; Boutron2010CONSORT; CARE_2013. Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian2021PRISMA; ISO29148_2018; IFRS_Integrated.
Narration & Style
Report Form corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian2021PRISMA; Boutron2010CONSORT; ICMJE2025; Leipzig_Rules; SIAM_Style; IPA_Handbook; APSA_Style; ASA_Style. Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron2010CONSORT; ICMJE2025; Chicago_Style; Blacks_Law; HBR_Guide. Paragraph reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian2021PRISMA; ISO29148_2018; The_Bluebook; ASA_Style. Reader Friendliness corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian2021PRISMA; ICMJE2025; HBR_Guide; GRI_Standards.
Ethics & Compliance
Sensitive Issues corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards ICMJE2025; Bastian2021PRISMA; OHA_Principles; OAH_Standards; BPS_Code; Belmont_Report; APSA_Ethics. Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication ICMJE2025; Guyatt2013GRADE; NRC_Prudent; Helsinki_Decl; ESA_Code. Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks ICMJE2025; Bastian2021PRISMA; BPA_GoodPractice; AAP_Code; CFA_Code; ABA_Model_Rules; AAPOR_Code.
Information Sufficiency
Source Support aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian2021PRISMA; FORCE11_Data; NBER2024; ISSCR2025; GIPS_2020. Information corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian2021PRISMA; ISO25010; EQUATOR_Network; LSA_Ethics2024; JOC_Guidelines; USGS_FSP; GRI_Standards. Citations align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability ICMJE2025; Wilkinson2016FAIR; IEEE2025Ethics; ACS2021Ethical; The_Bluebook. References are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base StanfordAI2025; FORCE11_Data; DA_RT_Principles.
Information Integrity
Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson2016FAIR; APS2023Conduct; NIST_AI_RMF; PDG_Review; NIST_DLMF; SEC_Reg_SK. Citation Validity corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson2016FAIR; ASA_Ethics; AHA2024Standards; Austin_Principles. Reference Accuracy aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson2016FAIR; NBER2024; FORCE11_Data; APA_JARS2025; IFRS_Standards. Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials ACS2021Ethical; IEEE2025Ethics; SEP_Policy; Helsinki_Decl; Belmont_Report. Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR_Network; LSA_Ethics2024; AAPOR_Code.
To establish a robust and universally applicable Deep Research Report Evaluation Taxonomy, this framework comprehensively integrates authoritative standards and guidelines from a wide spectrum of disciplines, ranging from quantitative sciences and engineering to the humanities, social sciences, and professional fields such as law and finance (see Table 10). By synthesizing the core principles shared across these diverse domains, this taxonomy prioritizes criteria that are universally recognized as essential for high-quality intellectual work. Consequently, the resulting evaluation model ensures that the generated reports not only adhere to domain-specific best practices but also satisfy the fundamental requirements of validity, clarity, and professional integrity demanded by the global research and practitioner communities.
Domain (80 listings) Reference
Artificial Intelligence (8) NIST_AI_RMF; StanfordAI2025; ACL2024Guidelines; NeurIPS2025Guidelines; ICML2025Instructions; ICLR2024Guidelines; JMLR2024AuthorGuide; Springer2025MLJournal
Biology (4) Bastian2021PRISMA; Boutron2010CONSORT; ISSCR2025; Percie2020ARRIVE
Business (4) GRI_Standards; HBR_Guide; ISO30414; IFRS_Integrated
Chemistry (4) ACS2021Ethical; IUPAC_Green; JOC_Guidelines; NRC_Prudent
Computer Science (4) ISO25010; ACM_Badging; ISO29148_2018; Wilkinson2016FAIR
Earth & Env. Science (4) IPCC_2019; USGS_FSP; CF_Conventions; ESA_Code
Economics (4) NBER2024; AEA2024Data; Econometric_Soc; OECD_Quality
Education (4) WWC_Handbook; AERA2006Standards; IES_SEER; AERA_APA_NCME_Standards
Engineering (4) IEEE_Std2025; IEEE2025Ethics; NASA_SE_Handbook; ISO29148_2018
Finance (4) IFRS_Standards; CFA_Code; GIPS_2020; SEC_Reg_SK
History (4) AHA2024Standards; Chicago_Style; OHA_Principles; OAH_Standards
Law (4) The_Bluebook; Blacks_Law; Fed_Rules_Evidence; ABA_Model_Rules
Linguistics (4) LSA_Ethics2024; Leipzig_Rules; Austin_Principles; IPA_Handbook
Mathematics (4) EMS_Code2025; AMS_Handbook; SIAM_Style; NIST_DLMF
Medicine (4) Helsinki_Decl; STROBE_2007; CARE_2013; SPIRIT_2013
Philosophy (4) APA_GoodPractices; SEP_Policy; BPA_GoodPractice; AAP_Code
Physics (4) APS2023Conduct; IUPAP_Red; PDG_Review; CODATA_2022
Political Science (4) DA_RT_Principles; APSA_Style; OECD_Policy_Eval; APSA_Ethics
Psychology (4) APA_JARS2025; OSF_Prereg; BPS_Code; AERA_APA_NCME_Standards
Sociology (4) AAPOR_Code; COREQ_2007; Belmont_Report; ASA_Style
Table 10: References by Domain used in Deep Research Report Evaluation Taxonomy. (Total 80 listings from 78 unique sources; 2 interdisciplinary standards are cross-listed).
Appendix E Rubric Structure
In this appendix, we summarize the hierarchical structure and scoring method of the rubric used to evaluate the Deep Research report.
Dimension / Criterion Element / Rubric description
2. Evidence Validity └ 2.2 Logical Support 2.2.5 Element All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps. [2.2.5.1 Coverage] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the Expert Evaluation Guidance, EG), and no central claim is left without the necessary supporting evidence. [2.2.5.2 Quality] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Evidence Validity / 2.2 Logical Support.
E.1 Hierarchical Levels
We have the following 4-level hierarchical structure to systematically evaluate report quality in multiple dimensions.
Level 1: Evaluation Dimensions (7)
The 7 upper dimensions presented in Table 1 represent different evaluation perspectives of report quality.
Level 2: Detailed Criteria (26)
Detailed criteria that decompose each dimension into finer evaluation units; 2–5 detailed criteria are defined under each dimension. This forms an intermediate-level unit that specifies which aspects to examine for each dimension.
Level 3: Evaluation Elements (40)
Evaluation elements that concretize each detailed criterion into units actually observable in the report. One or more elements are mapped under one detailed criterion, and are defined as evaluation units directly corresponding to report content, such as "Inclusion of requested items," "Specification of scope/limitations," and "Specificity of final message."
Level 4: Rubric Factors (130)
Rubric factors corresponding to the fourth level of the rubric are defined to evaluate each evaluation element (Level 3) from two perspectives. One element is subdivided into multiple factors from the perspectives of Coverage (C) and Quality (Q), and these factors serve as the minimum unit for actual scoring.
Thus, with 7 dimensions, 26 detailed criteria, and 40 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 130 specific check items. Table 11 shows an example of one element and its corresponding Coverage/Quality factors in the 2. Evidence Validity / 2.2 Logical Support dimension.
E.2 Coverage and Quality Factors
The rubric factors corresponding to Level 4 of the rubric evaluate each evaluation element (Level 3) from two perspectives: Coverage(C) and Quality(Q). The Coverage factor examines whether the element has been covered without omission as required throughout the document, i.e., how faithfully it includes content scattered in multiple locations or composed of multiple detailed requirements. On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., "how well it is written." By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report. Both factors use a 1–10 point scale, and detailed criteria for scoring are presented in Table 12.
Score Range Coverage Quality
9–10 (Perfect) Fully meets all requirements; No omissions; No revision needed Excellent quality in all relevant aspects; No revision needed — Top-tier international journal level, or high-end professional report meeting or exceeding standards in specific technical/industrial contexts
7–8 (Excellent) Meets almost all requirements; Only 1–2 minor omissions, minimal impact High quality; Meets most academic/professional standards, only minor improvements possible — Solid peer-reviewed journal, excellent doctoral research, high-quality industry report level
5–6 (Good) Meets more than half; Meets most key requirements, minor elements missing Meets essential professional standards; Clear structure and competent analysis but room for improvement — Well-written master’s thesis or standard professional report level
3–4 (Inadequate) Partially meets; Several key omissions Noticeable flaws in several aspects; Significant revision needed — Undergraduate thesis or entry-level professional report level
1–2 (Poor) Most requirements are missing or treated only superficially Fails to meet basic professional standards; Lacks depth, rigor, accuracy — Below undergraduate level; Unsuitable for publication or professional use
Table 12: Interpretation of 1–10 score ranges for Coverage (C) and Quality (Q) factors.
E.3 Scoring and Aggregation
The score aggregation flow in this evaluation framework consists of factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension. First, rubric factors (the lowest items) are evaluated with a 1–10 score (or N/A). Coverage (C) and Quality (Q) are calculated separately up to the factor/element level, then integrated at the element level, and the resulting element score is raised to the criterion and dimension levels. N/A items are excluded from all average calculations.
Let $r$ be the report, and type $T\in\{C,Q\}$ represent Coverage and Quality, respectively. Let the score of each rubric factor $f$ be $s_{r,f}\in\{1,\dots,10\}$ (or N/A), and the set of non-N/A factors of type $T$ belonging to element $e$ be $F_{e,r}^{T}$. Let $E_{c,r}$ be the set of elements belonging to criterion $c$, and $\mathcal{C}_{d,r}$ be the set of criteria belonging to dimension $d$. Then, the factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension aggregation is defined as follows:
The Coverage/Quality score (element level) of report $r$ for element $e$ is
$S_{r}^{T}(e)=\frac{1}{\lvert F_{e,r}^{T}\rvert}\sum_{f\in F_{e,r}^{T}}s_{r,f},\quad T\in\{C,Q\},$ (3)
and if $F_{e,r}^{T}=\varnothing$, $S_{r}^{T}(e)$ is set to N/A.
The integrated element score $S_{r}(e)$ of element $e$ is set as the average of defined values among C/Q. Let the set of types defined in element $e$ be $\mathcal{T}_{e}=\{T\in\{C,Q\}\mid S_{r}^{T}(e)\neq\text{N/A}\}$.
$S_{r}(e)=\frac{1}{\lvert\mathcal{T}_{e}\rvert}\sum_{T\in\mathcal{T}_{e}}S_{r}^{T}(e),$ (4)
If $\mathcal{T}_{e}=\varnothing$, $S_{r}(e)$ is set to N/A.
The score of criterion $c$ is defined as the average of element scores belonging to that criterion.
$S_{r}(c)=\frac{1}{\lvert E_{c,r}\rvert}\sum_{e\in E_{c,r}}S_{r}(e).$ (5)
If $E_{c,r}=\varnothing$, $S_{r}(c)$ is set to N/A.
The score of dimension $d$ is defined as the average of criterion scores belonging to that dimension.
$S_{r}(d)=\frac{1}{\lvert\mathcal{C}_{d,r}\rvert}\sum_{c\in\mathcal{C}_{d,r}}S_{r}(c).$ (6)
If $\mathcal{C}_{d,r}=\varnothing$, $S_{r}(d)$ is set to N/A.
In summary, we calculate $S_{r}^{C}(e),S_{r}^{Q}(e)$ by averaging C/Q factors in each element, integrate them to create the element score $S_{r}(e)$, then average the element scores by criterion to calculate $S_{r}(c)$, and finally average the criterion scores to calculate the dimension score $S_{r}(d)$.
Appendix F Human-based Information Verification Protocol
F.1 Overview
DREX’s Information Verification Protocol is based on a stepwise information verification procedure performed by human evaluators. This section details the actual two-step procedure (Claim Extraction, Factual Accuracy Evaluation) performed by human evaluators.
F.2 Step 1: Claim Extraction and Classification
Human procedure.
Evaluators segmented the report into sentences (in ‘Lx.Sy‘ format), reviewed each sentence individually to identify sentences containing claims, and extracted only the core claims. Pronouns were replaced with explicit references according to the context, and if a single sentence contained multiple claims, they were separated. All claims were classified into 6 types (A–F): Explicit Citation (A), Implicit – Same Section (B), Implicit – Previous Section (C), Structural Recap (D), No Citation Required (E), and Unknown Source (F). For A–C type claims, the corresponding citation or evidence position was recorded together. This process was performed on 2 randomly selected reports (total 728 claims), thereby constructing a Ground Truth for evaluating the recall of the extraction model.
Table 13 shows the definitions of the 6 claim types, and Table 14 shows examples of extracted sentences and classification results.
Table 13: Information Verification ModuleA–F Claim Type Definitions and Examples. Types A–C are targets for external evidence verification, while Types D–E–F are classified as internal information or unverifiable areas.
Type Definition and Example
A Cited Claim: A claim explicitly including a citation marker within the sentence. Example: “Multi-junction solar cells achieve efficiencies above 45% [1].”
B Uncited – Same Section / Paragraph: When the evidence citation exists in a previous sentence within the same section (or paragraph). Example: “This efficiency improvement is due to the layered structure.” (Evidence → L2.S1)
C Uncited – Previous Section / Paragraph: When the citation evidence exists in a previous section or paragraph. Example: “These findings confirm the results of earlier solar-cell studies.” (Evidence → L1.S3)
D Uncited – Structural Recap: A claim corresponding to a restatement of content in the document structure, such as introduction, conclusion, or summary. Example: “In conclusion, this paper reviewed recent advances in solar technology.”
E Uncited – No Citation Required: The author’s direct results, general knowledge, or a claim not requiring citation. Example: “Photosynthesis converts light energy into chemical energy.”
F No Citation – Unknown Source: A claim requiring external evidence but for which no source is presented. Example: “These panels can last for 50 years without degradation.”
Table 14: Example of human-annotated claims and corresponding LLM classification results.
Pos. Sentence / Extracted Claim Class (label)
L1.S3 The development of multi-junction solar cells has achieved efficiencies above 45% in laboratories [1]. A (Explicit citation)
L2.S1 This dramatic increase is due to the layering of different semiconductor materials. B (Implicit citation → linked to L1.S3)
L2.S3 The enhanced efficiency of these cells will reduce the land area required for solar farms. C (Implicit cross-section)
L2.S4 These new panels are durable enough to withstand a Category 4 hurricane. F (No known source)
F.3 Step 2: Claim Verification
Human procedure.
For 100 claims randomly selected from types A–C, two human evaluators independently assessed their factuality. The evaluation followed a single-criterion protocol determining whether the cited source explicitly supports (Supported) or lacks information/is irrelevant to (Not Supported) the content of the claim.
Annotator Qualifications & Adjudication.
The evaluators consisted of 2 individuals holding a master’s degree or equivalent experience in the report’s domain. The two evaluators made judgments independently, and for items where disagreement occurred, the final label unanimously agreed upon through discussion was established as the Ground Truth. Through this process, personal bias was excluded, and the objectivity of the evaluation was secured. The human verification results for these 100 claims were used as Ground Truth for the model performance evaluation in Section 6.5.
Detailed Verification Rubric.
Fact verification was strictly performed according to the following detailed criteria:
  • Supported: When the cited document clearly and directly includes the core facts (figures, causality, definitions, etc.) of the claim. The implied meaning in context must match the intent of the claim.
  • Not Supported: When the basis for the claim cannot be found in the cited document, or the document is irrelevant to the topic. Cases where the document link is invalid (404) or the content cannot be verified are also included.
Inter-human Agreement.
To verify the reliability of this protocol, we measured the agreement (Cohen’s Kappa) between the two evaluators. The analysis result showed that a high level of agreement (Substantial Agreement) of $\kappa=0.80$ was achieved in the Claim Support judgment. This suggests that the proposed verification criteria are objective and reproducible.
Appendix G LLM-based Information Verification Implementation
G.1 Overview
The Information Verification Moduleis designed to automate the human verification protocol described above. The LLM analyzes the report sentence by sentence to extract and classify claims, and if necessary, retrieves external documents to verify their factuality. In this process, algorithms such as Batch Extraction, Back-tracking, and Relevant Context Filtering were applied to achieve both cost efficiency and accuracy.
G.2 Claim Extraction and Classification
LLM adaptation.
The Information Verification Moduleautomates the procedure of human evaluators. After tokenizing the report into sentences, the LLM evaluates the "claim likelihood" of each sentence and predicts A–F types based on the 6-Type definitions defined in Table 13. For Type B and C claims, evidence sentences are subsequently searched using Back-tracking.
Batch Extraction Strategy
To mitigate the "Lost-in-the-Middle" phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy. After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$). Each batch processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences belonging to the current batch ($s_{i},\dots,s_{i+B-1}$).
Batch Extraction Prompt Structure System: You are an expert fact-checker and claim extractor. User: # Full Report Context {Full_Report_Text} # Target Sentences to Extract Claims From L10.S1: … L10.S20: … Extract claims only from the target sentences above. Use the full report context for coreference resolution.
Example JSON Output
 "claims": [{
  "position": "L10.S1",
  "index": 1,
  "claim_text": "Solar efficiency reached
  45% [1].",
  "claim_class": "A",
  "direct_citation": "[1]",
  "evidence_position": null
This method induces the model to focus on local sentences while remaining aware of the entire context, achieving human-level claim-extraction performance.
Classification Performance.
The results of performing type classification on 728 claims using GPT-5-mini are shown in Table 15. Explicit Citation (A) recorded a very high F1 score of 0.95, being identified almost perfectly. The distinction between Implicit Citation (BC) and Internal Information (DE) was somewhat challenging, but it recorded meaningful classification performance with F1 scores of 0.52 and 0.53, respectively. In particular, for the Unknown Source (F) type, which is the most difficult to identify, it achieved a precision of 0.35 and a recall of 0.25, indicating the ability to detect potential hallucination sentences. The overall accuracy was 0.68, which is a decent level considering the context of complex academic reports.
Table 15: Claim Classification Performance (GPT-5-mini). Classes are merged into A (Explicit), BC (Implicit), DE (Internal/Structural), and F (Unknown).
Class Precision Recall F1-Score Support
A (Explicit) 0.92 1.00 0.95 272
BC (Implicit) 0.42 0.68 0.52 163
DE (Internal) 0.76 0.41 0.53 242
F (Unknown) 0.35 0.25 0.30 51
Overall 0.68 0.68 0.67 728
Comparison with Sliding Window Baseline.
To validate the effectiveness of the LLM’s targeted evidence prediction, we compared it with a “Sliding Window” baseline on the subset of correctly classified B/C claims ($N=131$). The sliding window method collects all citations within a window of size $k$ centered on the claim. As shown in Table 16, the proposed LLM method achieves the highest Jaccard Index (0.7070) and Precision (0.7109), outperforming the sliding window baselines ($k=5,10,15$). While increasing the window size ($k$) improves Recall (up to 0.93), it significantly degrades Precision and Jaccard due to the inclusion of irrelevant citations. This result demonstrates that the LLM’s “Evidence Position” prediction provides a precise pointer to the supporting evidence, which is crucial for efficient verification.
Table 16: Baseline comparison on correctly predicted B/C claims ($N=131$). The GPT-5-mini method achieves the best balance of precision and accuracy (Jaccard).
Method Jaccard Precision Recall
Backtracking (Ours) 0.7070 0.7109 0.7383
Sliding ($k=5$) 0.6822 0.6822 0.8022
Sliding ($k=10$) 0.6247 0.6247 0.8791
Sliding ($k=15$) 0.5627 0.5627 0.9341
G.3 Claim Verification
Semantic Back-tracking for Citation Recovery
The Information Verification Moduleuses a Backtracking algorithm to find evidence for claims without explicit citations (Types B and C). Type B (Same Section) and Type C (Previous Section) claims often share citations from previous sentences in context. The algorithm traces the ‘evidence_position‘ (location ID of the reference target sentence, e.g., L1.S3) recorded in the claim’s metadata and adds the citations held by that target sentence to the citations of the current claim. This restores omitted citation relationships and allows the corresponding source to be reviewed together in the subsequent verification step.
Formally, for a set of claims $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$, each claim $c_{i}$ has a position $p_{i}$, type $t_{i}$, explicit citations $R_{i}$, and a reference position $ref_{i}$ (if $t_{i}\in\{B,C\}$). The Back-tracking function $f_{backtrack}(c_{i})$ is defined as follows:
$f_{backtrack}(c_{i})=\begin{cases}R_{j}&\text{if }t_{i}\in\{B,C\}\text{ and }\exists c_{j}\text{ s.t. }p_{j}=ref_{i}\\ \emptyset&\text{otherwise}\end{cases}$
Finally, the citation set used for verification becomes $R^{\prime}_{i}=R_{i}\cup f_{backtrack}(c_{i})$.
Relevant Context Filtering (Retrieval-Augmented Verification)
Using the entire report or long retrieved documents as input in the verification step is costly and can induce hallucinations due to unnecessary information (Noise). To solve this, we apply Relevant Context Filtering. For the verification target claim $q$ and the retrieved document $D_{retrieved}$, the document is divided into chunks $K=\{k_{1},k_{2},\dots,k_{r}\}$ (Size $\approx$ 1000 tokens). Then, the relevance score $Sim(q,k_{j})$ between each chunk and the claim is calculated, and only the top $N$ (Top-K) chunks are selected and used as input for the verification model.
In this study, the BM25 robertson2009probabilistic algorithm was used as the default Selector: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
$C_{final}=\text{Concatenate}(\text{SortByPosition}(K_{selected}))$
Through this process, token costs can be reduced by more than 80% while maintaining or improving verification accuracy.
LLM Verification Logic.
The Information Verification Module’s automatic verifier is prompt-engineered to determine whether the given context supports the claim, identical to the human protocol above. The LLM infers whether the claim’s core content and numerical information match the source to make a final judgment.
Example.
Claim: “Multi-junction solar cells achieve efficiencies above 45% in lab settings [1].”
Reference [1]: Reports a 46.2% lab efficiency under concentrated light.
Evaluation: The reference explicitly states 46.2% efficiency, supporting the claim of "above 45%".
$\Rightarrow$ Final Result: Supported
Fair Use and Direct Quote Detection
To ensure compliance with copyright and research ethics, the appropriateness of direct quotes is evaluated. First, the similarity between sentences in the report and the original text (Reference) is calculated using RapidFuzz and SequenceMatcher. If the similarity ($\tau$) is 0.9 or higher, it is considered a direct quote. Sentences detected as direct quotes are primarily checked for quotation marks (" ") or block quote formatting using regular expressions. If the format is insufficient or additional judgment is needed, the LLM is presented with the sentence and the original text to comprehensively evaluate compliance with "Fair Use Guidelines" (citation indication, minimal quotation, etc.). The final score is calculated by deducting points from a maximum of 10 based on the non-compliant ratio among all sentences.
G.4 Evaluation Metrics
The Information Verification Modulecalculates the following detailed metrics to evaluate Integrity and Sufficiency.
Integrity Metrics
  • External Claim Accuracy: The proportion of claims verified as factual among claims requiring external evidence (Type A, B, C).
[…]
  • Citation Accuracy: The proportion of citations that correctly support the corresponding claim among all citations.
    $\text{Cit. Acc}=\frac{|\text{Supported Citations}|}{|\text{Total Citations}|}$
  • Reference Accuracy (Supported per Shown): The proportion of references that actually contributed to content verification (Supported) among unique references shown in the report.
    $\text{Ref. Acc}=\frac{|\text{Supported Unique References}|}{|\text{Total Unique References Shown}|}$
  • Reference Reproducibility: The proportion of references that were accessible during the actual verification process and for which the webpage in markdown format could be successfully retrieved using the Jina API (not Error).
    $\text{Reproducibility}=1-\frac{|\text{Error References}|}{|\text{Used References}|}$
  • Information Reliability: The proportion of references that are both reliable sources and support the content among the used references.
    $\text{Reliability}=\frac{|\text{Reliable \& Supported References}|}{|\text{Used References}|}$
  • Reference Diversity (Citations CV): Uses the Coefficient of Variation of citation counts to measure dependency on specific references. (Lower is better; inverted for scoring)
    $\text{CV}=\frac{\sigma(\text{Citation Counts})}{\mu(\text{Citation Counts})}$
Sufficiency Metrics
  • Externally Verifiable Claims Ratio: The proportion of claims verifiable with external evidence (Type A, B, C) among all claims.
    $\text{Verifiable Ratio}=\frac{|\text{Claims (A, B, C)}|}{|\text{Total Claims}|}$
  • Information Quantity: The total number of claims verified as factual (Supported). (Raw Count before normalization)
    $\text{Info. Qty}=|\text{Accurate Verifiable Claims}|$
  • Citation Quantity: The total number of valid citations (Supported Citation) supporting claims.
    $\text{Cit. Qty}=|\text{Supported Citations}|$
  • Reference Quantity: The total number of valid references (Supported Reference) supporting claims.
    $\text{Ref. Qty}=|\text{Supported References}|$